OpenStudy (anonymous):

(3.6 times 10^5)(4.2 times 10^3) divided by (6 times 10^3)

OpenStudy (anonymous):

numerator = (3.6*4.2)*10^8 (powers can be added together when they have the same base) =15.12*10^8 now divide out the numerator =(15.12/4.2)*(10^8/10^3) ((powers can be subtracted when dividing if they have the same base) =2.52*10^5

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

CEEASU why did you divide 4.2 and not 6? 6 was the denominator. or 6 x 10^3 was

OpenStudy (anonymous):

wait nvm i gotcha!